ryry:

A store charges $8.95 for an 8 ounce jar of maple syrup. The price for a 10 ounce jar of the same brand is $12.00. Which jar is the better buy? What is the difference in the unit prices? If necessary, round to the nearest cent.

justus:

Welcome to Question Cove.

justus:

How are you doing today?

ryry:

Good.

justus:

That's nice to hear.

ryry:

How about you?

justus:

I'm doing just fine, thank you for asking.

justus:

I am here to help you.

justus:

Cost of 8 ounce jar is $8.95

justus:

Cost of 1 ounce jar is 8.95/8 =1.12

justus:

Cost of 10 ounce jar is $12.00

justus:

Cost of 1 ounce jar is 12/10=1.2

justus:

So 8 ounce jar is better since its unit price is less than 10 ounce jar. The difference in the unit prices=1.2-1.12=$0.08

ryry:

Thank you so much. This is very helpful

justus:

You're welcome. :)

justus:

Good luck.

ryry:

Thank you.

justus:

Thank you for the medal.

ryry:

You are very welcome.

justus:

:)

ryry:

:)